# Privacy-Preserved Neural Graph Databases

### v4

@@ 6.2. Benchmark Construction
Following previous work (Hamilton et al., 2018; Bai et al., 2023a; Wang et al., 2021b), we evaluate the complex query answering performance on the following eight general query types with abbreviations of $1p,2p,2i,3i,ip,pi,2u$, and $up$, as shown in Figure 4. For each general query type, each edge represents either a projection or a logical operator, and each node represents either a set of entities or numerical values. We use the sampling method proposed in (Bai et al., 2023a) to randomly sample complex queries from knowledge graphs. We random sample training, validation, and testing queries from the formerly constructed graphs respectively. For the training queries, we search for corresponding training answers on the training graph. For the validation queries, we use those queries that have different answers on the validation graph from answers on the training graph to evaluate the generalization ability. For the testing queries, we conduct a graph search on the testing graph with private edges for testing answers and we split the answers into private set and public set according to the definition of privacy risk query answers discussed in Section 4. We conduct a statistical analysis of the number of privacy risk answers for different types of complex queries of three knowledge graphs and the statistics are shown in Table 2.
@@ 6.3.2. Evaluation Metrics
We first tune hyper-parameters on the validation queries for the base query encoding methods. We use the same parameters for NGDBs for a fair comparison. We tune the privacy penalty coefficient $\beta$ for three datasets respectively for utility privacy tradeoff for convenient illustration. For noise disturbance privacy protection, we adjust the noise strength to make the results comparable to NGDBs’ results (similar performance on public query answers).
@@ ACKNOWLEDGMENTS
The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20) and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the support from UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08).
@@ Appendix A Instantiations of Query Encoders
In this section, we present detailed formulas for parameterizing various logical operations utilized in this paper. Additionally, we employ the following three query encoders: GQE, Q2B, and Q2P, as the foundational components of our approaches. Furthermore, to effectively handle numerical attribute values, we integrate the Number Reasoning Network (NRN) (Bai et al., 2023a) with these query encoding methods.
@@ A.3. Particle Embeddings (Q2P)
Here, $\sigma$ and $\phi$ are the sigmoid and hyperbolic tangent functions, and $\odot$ is the Hadamard product. To allow information exchange among different particles, a scaled dot-product self-attention (vaswani2017attention) module is also incorporated,
@@
A.4. Number Reasoning Network (NRN)
To better deal with numerical values in the queries, we incorporated NRN into these query encoding processes. Here are the implementation details.
A.4.1. Relational Projection and Logic Operations on Entities
In the number reasoning network, established query encoding methods are adopted as backbones for conducting relational projections and logic operations on entities. Three different encoding structures are utilized to encode the entity phase: query2box (Ren et al., 2020), GQE (Hamilton et al., 2018), and query2particles (Bai et al., 2022). Despite their differences, the query embeddings can be uniformly flattened to vectors, resulting in query encodings represented as $d$-dimensional vectors: $q_{i}\in\mathbb{R}^{d}$.
A.4.2. Attribute Projection
In the attribute projection step, given the query embedding $q_{i}$ of the entity variable at the $i$-th step and the attribute type $a$, we employ a neural structure $F_{p}$ to parameterize the distribution parameters for the $(i+1)$-th step:
(22) $\theta_{i+1}=F_{p}(q_{i},a),$
where $F_{p}$ is a neural structure whose weights are optimized during the learning process. A common parameterization for $F_{p}$ is the gated transition, inspired by the recurrent neural network unit. The parameterization equations are as follows:
(23) $\begin{split}&p_{i}=W^{p}_{p}q_{i}+b_{p}^{p}\\ &z_{i}=\sigma(W^{p}_{z}e_{a}+U^{p}_{z}p_{i}+b_{z}^{p}),\\ &r_{i}=\sigma(W^{p}_{r}e_{a}+U^{p}_{r}p_{i}+b^{p}_{r}),\\ &t_{i}=\phi(W^{p}_{h}e_{a}+U_{h}(r_{i}\odot p_{i})+b^{p}_{h}),\\ &\theta_{i+1}=(1-z_{i})\odot p_{i}+z_{i}\odot t_{i}.\end{split}$
where $e_{a}\in\mathbb{R}^{k}$ is the embedding used to encode the attribute type $a$. Here, $\sigma$ and $\phi$ represent the sigmoid and hyperbolic tangent functions, respectively, and $\odot$ denotes the Hadamard product. The parameter matrices are $W^{p}_{p}\in\mathbb{R}^{d\times k}$ and $W^{p}_{z},W^{p}_{r},W^{p}_{h},U^{p}_{z},U^{p}_{r},U^{p}_{h}\in\mathbb{R}^{k\times k}$.
A.4.3. Reverse Attribute Projection
To find entities that possess a numerical attribute value from a given set of values, the reverse attribute projection operation is employed to compute the query embedding for the resulting entity set:
(24) $q_{i+1}=F_{r}(\theta_{i},a),$
where $F_{r}$ is a neural structure with weights that are optimized during the learning process. The gated transition function in Equation (23) can also be used for reverse attribute projection. However, for this operation, the weights are substituted with a different set of parameters: $W^{r}_{p}$, $W^{r}_{z}$, $W^{r}_{r}$, $W^{r}_{h}\in\mathbb{R}^{k\times d}$, and $U^{r}_{z}$, $U^{r}_{r}$, $U^{r}_{h}\in\mathbb{R}^{d\times d}$.
A.4.4. Numerical Projection
The numerical projections are directly applied to the distribution parameters. Since the numerical constraint of $f$ can be an arbitrary binary function from $\mathbb{R}^{2}$ to ${0,1}$, we represent the projection operation using the function $F_{f}$ as follows:
(25) $\theta_{i+1}=F_{f}(\theta_{i}).$
A.4.5. Intersection and Union on Attribute Values
Suppose we have $n$ attribute variables $\theta_{i}^{1},\theta_{i}^{2},...,\theta_{i}^{n}$ as inputs for intersection and union operations. Their intersection and union can be expressed as follows:
(26) $\begin{split}\theta_{i+1}=F_{I}(\theta_{i}^{1},\theta_{i}^{2},...,\theta_{i}^{n}),\\ \theta_{i+1}=F_{U}(\theta_{i}^{1},\theta_{i}^{2},...,\theta_{i}^{n}),\end{split}$
where $F_{I}$ and $F_{U}$ are the functions that compute the intersection and union, respectively.
For the functions $F_{I}$ and $F_{U}$, we utilize the DeepSet (zaheer2017deep) functions to parameterize them. Specifically, we employ a self-attention network followed by a feed-forward network to implement the permutation-invariant DeepSet function. Given the input $\theta_{i}=[\theta_{i}^{1},\theta_{i}^{2},...,\theta_{i}^{n}]$, their formulas are as follows:
(27) $a_{i+1}$ $=\texttt{Attn}(W_{q}\theta_{i}^{T},W_{k}\theta_{i}^{T},W_{v}\theta_{i}^{T})^{T},$
(28) $\theta_{i+1}$ $=\texttt{MLP}(a_{i+1}),$
where $W_{q},W_{k},W_{v}\in\mathbb{R}^{k\times k}$ are parameters used to model the input Query, Key, and Value for the attention module Attn.

### v5

@@ 6.2. Benchmark Construction
Following previous work (Hamilton et al., 2018; Bai et al., 2023a; Wang et al., 2021b), we evaluate the complex query answering performance on the following eight general query types with abbreviations of $1p,2p,2i,3i,ip,pi,2u$, and $up$, as shown in Figure 4. For each general query type, each edge represents either a projection or a logical operator, and each node represents either a set of entities or numerical values. We use the sampling method proposed in (Bai et al., 2023a) to randomly sample complex queries from knowledge graphs. We randomly sample training, validation, and testing queries from the formerly constructed graphs respectively. For the training queries, we search for corresponding training answers on the training graph. For the validation queries, we use those queries that have different answers on the validation graph from answers on the training graph to evaluate the generalization ability. For the testing queries, we conduct a graph search on the testing graph with private edges for testing answers and we split the answers into private set and public set according to the definition of privacy risk query answers discussed in Section 4. We conduct a statistical analysis of the number of privacy risk answers for different types of complex queries of three knowledge graphs and the statistics are shown in Table 2.
@@ 6.3.2. Evaluation Metrics
We first tune hyper-parameters on the validation queries for the base query encoding methods. We use the same parameters for NGDBs for a fair comparison. We tune the privacy penalty coefficient $\beta$ for three datasets respectively for utility privacy tradeoff for convenient illustration. For noise disturbance privacy protection, we adjust the noise strength to make the results comparable to P-NGDBs’ results (similar performance on public query answers).
@@ ACKNOWLEDGMENTS
The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20) and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the support from UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21,
RMGS23CR05, RMGS23EG08).
@@ Appendix A Instantiations of Query Encoders
In this section, we present detailed formulas for parameterizing various logical operations utilized in this paper. Additionally, we employ the following three query encoders: GQE, Q2B, and Q2P, as the foundational components of our approaches.
@@ A.3. Particle Embeddings (Q2P)
Here, $\sigma$ and $\phi$ are the sigmoid and hyperbolic tangent functions, and $\odot$ is the Hadamard product. To allow information exchange among different particles, a scaled dot-product self-attention (Vaswani et al., 2017) module is also incorporated,